OpenStudy (anonymous):

The perimeter of a rectangle is twice the sum of its length and it's width. The perimter is 40 meters and its length is 2 meters more than twice its width. What is it's Length ?

OpenStudy (anonymous):

my guess was 14

OpenStudy (a1234):

2w + (2w + 2)(2) = 40 2w + 4w + 4 = 40 6w = 36 w = 6 12 + 2 = 14- yes

OpenStudy (anonymous):

thanks (:

OpenStudy (a1234):

You're welcome.

OpenStudy (anonymous):

medal please ?